1. The heat of combustion of benzoic acid (C6H5COOH) is -3226 kJ/mol. When 0.766 g of benzoic acid was burned in a bomb calorimeter, the temperature of the calorimeter (including its contents) rose from 20.75 °C to 25.32 °C. What is the heat capacity (calorimeter constant) of the calorimeter?
2. Use the References to access important values if needed for this question.

Calculate the ΔH° for the following reaction using the information below it.
C2H5OH(l) + O2(g) CH3CO2H(l) + H2O(l); ΔH° = ?
C2H5OH(l) + 1/2 O2(g) CH3CHO(l) + H2O(l); ΔH° = -175 kJ
CH3CHO(l) + 1/2 O2(g) CH3CO2H(l); ΔH° = -318 kJ
1
Expert's answer
2019-10-31T07:29:50-0400

1)К=3226·0.766/(122·(25.32-20.75))=4.43

2)Н=-175+(-318)=493